OpenStudy (anonymous):

What is the vertex of the graph of y = 3x2 + 2x + 1? A.-1/3, 0 B.-1/3, 2/3 C.1/3, 2 D.2/3, 11/3

OpenStudy (anonymous):

y=3(0)^2 +2(0)+1 y=0+0+1 y=1

OpenStudy (anonymous):

th vertex (turning point) is the point at which it reaches zero

OpenStudy (anonymous):

no, it does not seem to be one of the options

OpenStudy (anonymous):

the solution involves complex numbers, so there is no simple solution

OpenStudy (radar):

When the parabola is given in the quadratic form, such as yours is: You can find the x coordinate of the vertex by using -b/(2a) For this particular parabola after plugging in the values of a and b, the x value is -2/2*3 or -1/3. So the vertex is at (-1/3 , y) Now it is only a matter of plugging in the x value in the original formula and solving for y to get the y value.

OpenStudy (radar):

Solving for y, with x at -1/3 Y=3(-1/3)^2 + 2 (-1/3) + 1 y= 3/9 -2/3 + 1 y = 2/3 So the vertex is at (-1/3, 2/3) I do believe you may have an option that is satisfactory. @brittany8432

OpenStudy (radar):

The vertex is the point where the parabola reaches its maximum or minimum depending on whether it is opening up or opening down. It may or may not occur when y = 0 or when x = 0

OpenStudy (anonymous):

As in local max or min

OpenStudy (radar):

Yes, y=3x^2 + 2x + 1 y'=6x + 2 set to 0 6x+2 = 0 6x=-2 x=-2/6 = -1/3 I did not know if the poster knew calculus. @BPDlkeme234